Consider a EU consumer who buys both expensive and inexpensive running shoes imported from a non-member country. The expensive shoes cost 140 euros and the inexpensive shoes cost 35 euros (without the tariff). If the tariff on the inexpensive shoes is 16.9%, use an indifference curve-budget line analysis to show how imposing the tariff affects the bundle of shoes the consumer buys compared to what she/he would have bought in the absence of the tariff. Can you predict whether she/he will buy relatively more expensive running shoes after the tariff ? Why or why not ?
Answer #1

Given,

cost of expensive shoes ( including tariff ) =140 euros

cost of inexpensive shoes ( without tariff ) = 35 euros

given tariff of 16.9 %, tariff on inexpensive shoes = 5.915 euros

so cost to customer of inexpensive shoes after tariff = 35 + 5.915 =40.915.

The before and after tariff budget constraints are as follows:

A) before tariff = 140x + 35y =M ( where M= income, x= expensive shoes, y=inexpensive shoes )

B) after tariff = 140x + 40.915y =M

The effect of tariff using indifference curve analysis can be seen using the following diagram:

The imposition of a tariff shifts the budget constraint downwards from M/140---M/35 to M/140----M/40.915

This leads to shifting to lower IC

Now, 2 IC's have been drawn that depict the substitution effect and the income effect.

Substitution effect states that since one of the product is now expensive ,therefore the consumer with substitute that product in favor of another product.

Income effect states that since the consumer is now poorer in real terms the consumption of both the commodities will be changing.

Hence the total effect depends on the relative strengths of the income and the substitution effects which are different from consumer to consumer.

If substitution effect is greater than income effect then point A will be the final allocation.

If income effect is larger, then point B will be the final allocation.

Therefore we cannot predict with certainty that he/she will be buying expensive goods more or less after the tariff imposition.

Enpensine shoes m 1115 M 15 Thompensive Shoes